a uniform plane electromagnetic wave has wave length of6cm in free space and 4cm in dielectric for which u=4.7×10^-7 na^-2.determine the dielectric constan
1
Expert's answer
2020-04-02T10:44:09-0400

"\mu =4.7\cdot10^{-7}"


"\nu=\frac{c}{\lambda_0}=\frac{v}{\lambda}\to v=\frac{c\lambda}{\lambda_0}=\frac{3\cdot10^8\cdot0.04}{0.06}=2\cdot10^8m/s"



"v=\frac{c}{\sqrt{\mu \epsilon}}\to \epsilon=\frac{c^2}{v^2\mu}=\frac{(3\cdot10^8)^2}{(2\cdot10^8)^2\cdot4.7\cdot10^{-7}}=4.8\cdot10^7"